High connection latency, and general "things not working the way they are supposed to"

Wednesday, February 2, 2022 6:09 AM

Problem(s) Encountered:

The TCS responded VERY slowly to commands to open the dome. Then the mirror cover would not open at all.  I tried stopping and restarting the TCS, but the telescope server would not start.


...determined that the TCS power needed to be recycled in the control room.  Observer came down from the 2.4m and did that.  This causes the telescope focus to go to zero and the telescope coordinates have to be zeroed using the tilt meters. (I'm not sure why the telescope had to be moved in this case because I had not moved it from zenith yet.)  That took a while, but things have gone smoothly since.